OpenStudy (ray1998):

Please help :) fan and medal will be given. Thank you in advance! Left f(x) = 6x + 8 and g(x) = x^2. Perform the function operation and then find the domain of the result. (f + g)(x)

jimthompson5910 (jim_thompson5910):

`(f + g)(x)` is equivalent to saying `f(x) + g(x)`

jimthompson5910 (jim_thompson5910):

perform substitutions and rearrange terms to get it into standard form

OpenStudy (ray1998):

Okay, thank you! Can you show me how to set it up please? I'm sorry, I know I should know this, but I get confused.

jimthompson5910 (jim_thompson5910):

here's an example f(x) = 9x+17 g(x) = 3x^2 (f+g)(x) = f(x) + g(x) (f+g)(x) = 9x+17 + g(x) (f+g)(x) = 9x+17 + 3x^2 (f+g)(x) = 3x^2 + 9x+17 In step 1, I used the definition of what (f+g)(x) represents In step 2, I replaced f(x) with 9x+17, since they are equivalent In step 3, I replaced g(x) with 3x^2, since they are equivalent In step 4, I rearranged the terms to get the polynomial into standard form this is NOT the answer since it's just an example; however, this should hopefully help you out since it's a similar problem

OpenStudy (ray1998):

Thank you very much!!!! It does help :) I'm gonna look over it more and apply the steps to the actual problem I'm working on and see if I can get the right answer. I'm also going to go ahead and give you the best answer :) thank you for helping me!

OpenStudy (jdoe0001):

\(\bf f(x) = 6x + 8 \qquad g(x) = x^2 \\ \quad \\ (f+g)(x)\implies f(x)+g(x)\implies (6x + 8)\quad +\quad (x^2)\)

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

let me know if you're stuck on how to find the domain

OpenStudy (ray1998):

I will :) thank you, again

OpenStudy (ray1998):

Is it x2 + 6x + 17?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

idk where you got the 17 from

jimthompson5910 (jim_thompson5910):

the problem gives ` f(x) = 6x + 8 and g(x) = x^2` right?

OpenStudy (ray1998):

Yes! Sorry for the late reply.

jimthompson5910 (jim_thompson5910):

so it should be x^2+6x+8

OpenStudy (ray1998):

Yes, sorry! I meant 8. I got the example confused with the other problem.

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (ray1998):

And the domain is a set of all real numbers, correct? Or am I wrong?

jimthompson5910 (jim_thompson5910):

yep, the domain of f(x) is the set of all real numbers the domain of g(x) is the set of all real numbers the domain of (f+g)(x) is the set of all real numbers

jimthompson5910 (jim_thompson5910):

why? because each is a polynomial

OpenStudy (ray1998):

Awesome! Yay ^.^ thank you :) I have another problem, Let f(x) = x-2 and g(x) = 4x^2. Perform the function operation and then find the domain of the result. Would I apply the exact steps that you showed me from before? :)

jimthompson5910 (jim_thompson5910):

which operation? it doesn't say I'm guessing (f+g)(x) ? but again, it doesn't specify the operation

OpenStudy (ray1998):

Oh, sorry! I forgot to put that. No, the operation is (f-g)(x)

jimthompson5910 (jim_thompson5910):

we are given \[\Large {\color{red}{f(x)}} = {\color{red}{x-2}}\] \[\Large {\color{blue}{g(x)}} = {\color{blue}{4x^2}}\] and we want to find \[\Large (f-g)(x)\]

OpenStudy (ray1998):

Yes

jimthompson5910 (jim_thompson5910):

so... \[\Large (f-g)(x) = f(x) - g(x)\] \[\Large (f-g)(x) = {\color{red}{f(x)}} - {\color{blue}{g(x)}}\] \[\Large (f-g)(x) = ({\color{red}{f(x)}}) - ({\color{blue}{g(x)}})\] \[\Large (f-g)(x) = ({\color{red}{x-2}}) - ({\color{blue}{4x^2}})\] I'll let you finish up

OpenStudy (ray1998):

f = - 3x^2 - gx - x + 2 /x?

jimthompson5910 (jim_thompson5910):

huh? is this a different problem?